OpenStudy (anonymous):

help please , n/5 - 3n/10 = 1/5 what is n

OpenStudy (turingtest):

n=-2 care why?

OpenStudy (anonymous):

Make sure the bottom has a common denominator. So multiply the first term by 2/2 and the last term by 2/2. You should get: 2n/10 - 3n/10 = 2/10 Multiply it out by 10 to get 2n - 3n = 2 -n = 2 n = -2

OpenStudy (anonymous):

Multiply by a common factor, that is 10 to everything. You get 2n-3n=2. -n=2 so n=-2

OpenStudy (turingtest):

yup